OpenStudy (anonymous):

how can salve this first order ordinary differential equation by method separation variables dy/dx=1+y+x^2+yx^2

OpenStudy (anonymous):

Separable.

OpenStudy (anonymous):

\[ \frac{y'}{y+1}=1+x^2 \]

OpenStudy (anonymous):

\[ \int \frac{y'dx}{y+1}=\int 1+x^2\;dx \]\(dy=y'dx\)\[ \int \frac{dy}{y+1}=\int 1+x^2\;dx \]